OpenStudy (anonymous):

Damien wants to deposit $1200 into savings accounts at three different banks: Bank of Chance, Merchant Bank, and Utopian Financing. He will deposit three times as much into Merchant Bank as Bank of Chance because they offer a higher interest rate. He also expects the Utopian Financing deposit to be only 20% of the total of the other two deposits. How much will Damien deposit into the Utopian Financing savings account?

OpenStudy (anonymous):

$240 $250 $200 $750 Help please!

OpenStudy (anonymous):

M, C, and U are your amounts, and their sum is1200 3C=M .2(M+C)=U=4C/5 C+3C+4C/5=1200 Solve for C and substitute back into formula for M 4C+4C/5=1200 20C/5+4C/5=1200 24C=5*1200 C=250 U=4C/5=1000/5=200